OpenStudy (anonymous):

Find the value of x. 4.75 6.5 8 12

OpenStudy (anonymous):

okay but how do i do that provide me something

OpenStudy (anonymous):

ishaan has a point :O

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

you do

OpenStudy (amistre64):

yesterday it was 6.5; so im gonna presume since i can no longer assume that it is about 12 today

OpenStudy (amistre64):

with everything halved; id say that: 2x+5 = (6x-14)/2

OpenStudy (amistre64):

2x +5 = 3x -7 -2x +7 -2x +7 ---------------- 12 = x

OpenStudy (amistre64):

HAH!! i was right :)

OpenStudy (anonymous):

wat are yall talking about??